Project

General

Profile

Actions

Physics Shift Report

Dates: 09-10-2017 - 13-10-2017
Shifter: Christoper Hunt
Run Numbers: 09905 - 09965

Analysis Objectives

Table of Runs

Objective Analysis Code Channel Setting Run Numbers
lH2 Scattering S14d Field Off 3-240+M3-Test1 09935 09936 09937 09938 09939 09940 09941 09942 09943 09944 09945
lH2 Emittance Evolution H31a 2017-02-6 3-140+M3-Test3 09906
H31b 2017-02-6 4-140+M3-Test2 09905
H35a 2017-02-5 3-140+M3-Test3 09918 09920 09923 09926 9932
H35b 2017-02-5 6-140+M3-Test2 09921 09924 09927 09931
H35c 2017-02-5 10-140+M3-Test4 09922 09925 09929 09930
H36a 2017-02-7 3-140+M3-Test4 09947 09956 09961
H36aa 2017-02-7 3-140+M3-Test3 09952 09954 09955 09965
H36b 2017-02-7 4-140+M3-Test1 09948 09950 09957 09958 09962
H36c 2017-02-7 6-140+M3-Test2 09949 09959 09960 09963
H36d 2017-02-7 10-140+M3-Test4 09953 09953 09964
Momentum Scan/Alignment H32a 2017-02-6 3-170+M3-Test1 09911
H32b 2017-02-6 3-200+M3-Test2 09910 09915
H32c 2017-02-6 3-240+M3-Test1 09907 09908 09909 09912 09913 09914

Notes

Analysis Tasks: H33 & H34 were for beamline tuning and were scrutinised in depth by the owner. (C Hunt, J Pasternak, C Rogers, T Mohayai)

Analysis Task: H37 Another tuning run (30k Triggers) for the 4mm setting.

Elog, CDB, and Reconstruction

Elog entries: All ok.

CDB uploaded: All ok.

Online reconstruction:

Notes

CDB End Coil Currents

It was noted (whilst on shift) that the CDB was not receiving the correct current settings for the End coils from run control.

This was later rediscovered by C Rogers when trying to analyse emittance evolution - the field magnitude was wrong therefore the momentum reconstruction was wrong.

The CDB was retroactively moderated to correctly reflect the cooling channel settings, and the runs are being reprocessed. Hopefully the issue is resolved.

General

  • Noticed that the EMR normalised Chi-squared values appear to have a momentum dependence. Higher momentum runs have slightly larger chi-squared values. Still with expected limits however. May be due to relative pion contamination.
  • There is a PMT power supply issue in the EMR plane 26. This is known about by the experts so no problem.
  • Some runs have no OnRec plots on reco.mice at time of writing: 09927 09928
  • Noticed during the Field-Off runs that some demonstrate what appears to be a shadow of the diffuser in TKU-station5. Probably completely fine - just interesting.

Updated by Hunt, Christopher over 5 years ago · 23 revisions